OpenStudy (anonymous):

what is true about the complex number -5-5i a. the modulus is 5 sqrt2 b. it lies in quadrant 2 c. the modulus is 5 d. it lies in quadrant 3

OpenStudy (anonymous):

Think of that number as just a vector in the complex plane. If your complex number is a+bi modulus = sqrt(a^2 + b^2) To find the quadrant, draw it and check.

OpenStudy (anonymous):

i think it lies in quadrant 3 and not quadrant 2

OpenStudy (anonymous):

for the modulus is it 5?

OpenStudy (anonymous):

It is in quadrant 3 but modulus is not 5 modulus = sqrt(a^2+b^2)=sqrt((-5)^2+(-5)^2)=sqrt(25+25)= sqrt(50) = sqrt(2*25)=5sqrt(2)

OpenStudy (anonymous):

oh I see my mistake. Thank you :)